OpenStudy (waheguru):

How Do I convert 3pi/4 to degrees

OpenStudy (anonymous):

angle in degrees = angle in radians * 180 / Pi angle in degrees = 3 Pi / 4 radians * 180 / Pi angle in degrees = 3 / 4 * 180 angle in degrees = 135

OpenStudy (anonymous):

(3Pi/4)/2Pi=x/360

hero (hero):

Replace pi with 180 to get \(\frac{3(180)}{4}\) Then continue from there